The profit equation for John Billiard Supply company is given as follows: Profit=200X-1000-100X Find selling price...

  1. The profit equation for John Billiard Supply company is given as follows:

Profit=200X-1000-100XExjIap1Z2W6oe79LBFqufLgcM7fqrczcBVaqqIM2

  1. Find selling price per unit, fixed cost and variable cost per unit.
  2. If company sells 10 units, then find total expenses of the company?
  3. If company sells 15 units, then what will be the total profit of the company?

2. Ina Production Company, the materials and labor cost for making a product is $200 and the fixed cost per week is $2000. The selling price for each product is $300.

  1. If company sells 200 units, find total revenue for the company.
  2. How many products must be sold by the company per week to breakeven point?

3. A manager is deciding whether or not to build a small facility. Demand is uncertain and can be either at a high or low level. If the manager chooses a small facility and demand is low, the payoff is $30. If the manager chooses a small facility and demand is high, the payoff is $10. On the other hand, if the manager chooses a large facility and demand is low, the payoff is -$20, but if demand is high, the payoff is $80.

  1. Develop the decision table.
  2. What would be the best decision based on the maximax criterion?
  3. What would be the best decision based on Hurwicz's criterion of realism using α = 0.6?

*Answer these questions

*use MS format not Pic

* course is Quantitative Methods
Answer #1

(1)

Profit = TR - TC = 200X - 1000 - 100X, where

Revenue (TR) = 200X

Total cost (TC) = 1000 + 100X

(a)

Since TR = Price x Quantity,

Selling price per unit = 200

Fixed cost (FC) = 1000

Fixed cost per unit = FC/X = 1000/X

Variable cost = 100X

Variable cost per unit = VC/X = 100X/X = 100

(b)

When X = 10,

TC = 1000 + (100 x 10) = 1000 + 1000 = 2000

(c)

When X = 15,

Profit = (200 x 15) - 1000 - (100 x 15) = 3000 - 1000 - 1500 = 500

(2)

(a)

TR = Price x Quantity = $300 x 200 = $60,000

(b)

Break-even units = Fixed cost / (Unit price - Unit variable cost)

= $2000 / $(300 - 200)

= $2000 / $100

= 20 units

NOTE: As per Answering Policy, 1st 2 questions are answered.